# Towards Variable-Length In-Network Caching

### v2

@@ 3.2 Packet Format
Figure 3 depicts the StarCache packet format. The switch only parses the header. The payload of StarCache message consists of the key and the value. Our header is 22 bytes. Therefore, StarCache supports a key-value pair of up to 1438 bytes for a single packet. For example, the switch can cache an item with a 16-byte key and a 1422-byte value. Our header fields are as follows.
@@
Figure 4: Request processing. (a) the switch drops the request after inserting request metadata into the request table; (b) If a circulating cache packet reads request metadata, the switch clones the packet so that the original packet is forwarded to the client and the cloned one is recirculated again for further serving; (c) the switch invalidates the item to avoid inconsistent reads if a write request is for a cached item; (d) upon receiving a write reply for a cached item, the switch validates the item. After that, the switch clones the packet similar to (b). The cloned packet is recirculated as a read reply after the first recirculation.
@@
  • FLAG (1 byte): a versatile flag field 1) to distinguish write requests for cached items from those for uncached items; 2) to store the number of packets comprising a multi-packet item.
@@ 3.3 Basic Request and Reply Processing
Request generation at clients. To send a request, clients should specify the OP and HKEY fields. The SEQ field is increased by one for every request, which is used to resolve hash collisions. As we consider hash-partitioned storage, the destination storage server is determined by hashing the key.
@@
Read replies. When receiving a read reply, the switch first checks to see if the ingress port is the recirculation port. If it is, the reply is a cache packet; otherwise, it is a reply for an uncached item sent by the server. For cache packets, if the cache misses or the state of the value is invalid, the switch drops the packet. A cache miss for a cache packet means that the controller evicted the key from the lookup table due to a change in key popularity or the cache size. The invalid state indicates a write request with a new value is in progress.
The switch goes through the lookup table and the state table, and then looks for pending requests in the request table. If a request is found, the switch forwards the cache packet to the client after updating the header with metadata and removing the metadata from the table slot. Here, to make the cache packet serve more requests, the switch clones the packet before packet forwarding. The switch forwards the original packet to the client and the cloned one to the recirculation port. The switch also recirculates the packet if there are no pending requests in the table.
Write requests. The first switch checks whether the requested key is in the cache lookup table. If it is, the value for the key is invalidated to prevent reading the outdated value. In addition, the switch sets the FLAG field to 1 so that the storage server can append the value in the write reply. Regardless of a cache hit, the switch forwards the request to the storage server to update the value of the key in the server.
Write replies. If the key is cached, the switch validates the value to allow read requests to get the latest value. Similar to reads, the switch clones the packet so that the client receives the write reply while the switch has a new cache packet simultaneously. The switch updates the OP field to R-REP of the cloned packet after the first recirculation since cache packets are read replies. For a cache miss, the switch forwards the packet to the client since the reply is for an uncached item.
Other types of messages. Fetch requests, fetch replies, and correction requests are delivered to the server as standard packets. The fetch reply is forwarded to the controller, who has sent a fetch request for cache updates.
@@ 3.4 Buffering Request Metadata
Circular queue-based request table. We should store multiple request metadata for cached items because the requests should wait until cache packets serve them. The switch should be able to buffer concurrent read requests for the same cached item. If not, many requests for the cached item would be forwarded to the storage server due to the lack of vacancies. Furthermore, a request for a key should be isolated from other requests for other keys since the request may undergo excessive queueing delay due to the different requests.
We design a request table based on a circular queue structure. Our request table provides a logical queue for each cached key and the queue can be accessed in $O(1)$ by efficient indexing. This provides fast and isolated queue access for a key regardless of other keys. The request table consists of 6 register arrays: the first three arrays are for storing request metadata for each key (i.e., client IP address, request sequence number array, and L4 port number). The other three arrays are for managing queue operations. The queue length array maintains the number of stored request metadata for each key. The front pointer array handles dequeue operations by tracking the first request metadata for each key. The rear pointer arrays perform enqueue operations by keeping track of the last request metadata for each key.
The queue management arrays are indexed using a table index ($CacheIdx$) returned by the cache lookup table. The request metadata arrays are accessed via a request index, which is defined as $ReqIdx=CacheIdx\times S+i$ where $S$ is the maximum queue size per key and $i$ is the offset for $i$th slot in the logical queue for a key and is given by the pointer arrays. The switch uses three match-action stages for a request table. Specifically, the switch first checks the queue status (Stage 1), and performs en/dequeue operations if the queue is not empty or not full (Stage 2). Next, the switch gets/puts metadata (Stage 3).
@@ 3.5 Cache Serving
Handling multiple requests via packet cloning. In StarCache, cache packets pass through the switch data plane repeatably to serve pending requests for cached items. One challenge is serving multiple requests with a cache packet fetched only once because the cache packet is forwarded to the client after updating the header using the retrieved metadata. A strawman is to fetch the cache packet from the server again, but this approach is inefficient as the switch cannot serve pending requests for the key until the fetching is completed. Fetching multiple copies of the cache packet may be a solution, but this incurs excessive queueing delay between cache packets. In addition, it is hard to know how many cache packets should be fetched because we cannot predict the number of requests for each cached item in advance.
@@
To address this, we utilize packet cloning, another built-in feature of the programmable switch ASIC besides packet recirculation. Packet cloning is done by the PRE, a hardware module specialized for packet cloning in the switch ASIC. Cloning has a negligible overhead for the following reasons. First, the PRE is located after the ingress pipeline. This means the switch does not repeat the ingress pipeline processing for the cloned packet, not causing extra ingress processing delay. Second, the switch does not copy the entire packet. It only copies the small descriptor pointing to the memory location of the packet and reuses the packet.
@@ 3.6 Handling Hash Collisions
Client-side collision resolution. We handle hash collisions at the client level by maintaining a list of the keys for each request that has not yet received a reply. The list is indexed by $pkt.seq$, a request ID. Request packets contain both the original key and the key hash since we should use the original key to get the value in servers. Upon receiving read replies, the client checks whether the requested key in the list and the returned key in the reply header are identical. If different, the client sends a correction request to the storage server. The switch then forwards the request without applying the cache logic so that the client gets the correct item value from the storage server. One issue is that requests for uncached but hash-collided keys always undergo this process. However, this is uncommon since our 128-bit key hash provides a low collision probability. In our experience, hash collisions only occur occasionally during cache updates.
An example of resolving hash collisions. We plot an example of handling hash collisions in Figure 7 where the requested key is DDDD but the returned key is AAAA. Upon receiving the reply, the client detects that the retrieved value is a wrong value by referring to the key list using $pkt.seq=1$. The client then sends a correction request to the storage server. The switch bypasses the cache logic, and forwards the packet to the server. The storage server returns the value as a read reply, and the client finally gets the correct value for the key DDDD. The client removes the key from the list. $pkt.seq$ is finite and wraps around if it reaches the maximum value.
@@ 3.7 Cache Coherence
To ensure cache coherence between the switch and servers, we design an invalidation-based coherence protocol illustrated in Figure 4 (c) and (d). The switch invalidates the value when handling a write request for a cached item and revalidates the value upon receiving a write reply. Ideally, the outdated cache packets should be replaced with the latest cache packet. However, requests may be returned with the stale value since the outdated packets are still circulating in the switch data plane even if the item value becomes invalid. To address this, the switch drops the cache packet if the item is cached but its value is invalid. Owing to the fact that the cache packet is dropped before accessing the request table, we can prevent read requests from retrieving the stale value until a new cache packet is fetched.
In StarCache, the server sends a single reply packet, and the switch updates the value and replies to the client simultaneously by cloning the packet. Note that the existing architecture [22] sends two different packets for updating the value and replying to the client, which degrades write performance.
@@ 3.8 Cache Updates
StarCache can handle dynamic workloads by tracking workload changes as follows.
Dealing with key popularity changes. To handle key popularity changes over time, we update cache entries based on periodic popularity reports from the switch and storage servers. In particular, we count the popularity of cached keys in the switch data plane using the key popularity counter. The controller keeps track of key popularity by reading the counter periodically. Meanwhile, storage servers periodically report the top-$k$ keys to the controller, which are popular uncached keys. The servers use a count-min sketch with five hash functions to track key popularity in a memory-efficient manner while ensuring accuracy. To reflect the recent status only, we reset all the counters to zero after reporting.
If cache entries need to be updated due to a change in key popularity, the controller removes the victim key from the cache lookup table and inserts a new popular key instead. The controller then sends a fetch request to the storage server containing the latest value for the key. Upon receiving the request, the server fetches a new cache packet to the switch data plane and replies to the controller.
Dynamic cache sizing. Since StarCache uses recirculation, it is important to avoid the queueing between cache packets. If not, many requests for cached items may be forwarded to the storage server since they did not find a free slot in the request table where cache serving is delayed due to the queueing between cache packets. Although we try to set the best cache size at first, the cache size may cause a negative effect as the workload changes over time.
To address this, the controller dynamically changes the cache size to maximize the performance while avoiding the adversarial effect. The controller calculates the request overflow ratio, which is defined as the number of requests for cached items that were sent to the server divided by the total number of requests for cached items. The controller compares the ratio with a pre-defined threshold (e.g., 1%). If the ratio exceeds the threshold, the controller decreases the cache size by one. Otherwise, the cache size is increased by one.
@@ 3.9 Handling Practical Requirements
Multi-packet items. Many key-value items are less than the MTU size. However, some items may exceed the MTU size, such as articles and photo objects (e.g., Wikipedia and Flickr workloads [9]). To cache multi-packet items, we fetch multiple cache packets with fragmented values for the same key. We maintain the number of forwarded cache packets for each item by placing another register array alongside the request table, the ACKed packet counter. The initial value of each slot is 1 since most items are single-packet. When fetching item values, the storage server puts the number of packets comprising the item in the FLAG field. When handling cache packets, the switch retrieves metadata as usual, but it does not manipulate the slots of all register arrays in the request table if the slot value in the ACKed packet counter is not equal to the value in FLAG. Instead, the switch increases the slot value by one. When the slot value in the counter equals FLAG, the switch removes request metadata as it is ready to be finished.
However, this may reduce the performance gains since StarCache has to limit the cache size to avoid queueing between cache packets, and multiple cache packets occupy the cache for one large item. Therefore, the balancing efficiency may decrease if many items consist of multiple packets. This can be mitigated if the bandwidth of the recirculation port increases in next-generation switch hardware.
Failure handling. For packet loss, we can use an application-level mechanism. For example, our controller uses UDP with a timeout-based mechanism to send fetch requests and receive fetch replies. Storage servers use TCP to report top-$k$ items to the controller. Meanwhile, server failures do not cause a StarCache-specific problem. Switch failures result in the loss of cached items, but the cache can be reconstructed quickly after the switch is recovered.
Multi-rack deployment. StarCache can scale to multiple racks as items of servers are only cached in the ToR switch. For example, assume that we have rack 1 and rack 2 where ToR switches, $ToR1$ and $ToR2$, are interconnected via a spine switch $SPN$. When $CLI$, a client in rack 1, sends a request to $SRV$, a server in rack 2, the path is $CLI-ToR1-SPN-ToR2-SRV-ToR2-SPN-ToR1-CLI$. If the requested item is cached, $ToR2$ is the only switch that applies the cache logic. There is no issue even if rack 1 is the replicated rack of rack 2 where the two ToR switches have the same cache entries because the item is returned by $ToR1$.
@@ 3.10 Discussion
Write-back caching. Although most key-value workloads are read-intensive, some workloads are write-intensive. Since we use write-through caching that updates the switch and the server simultaneously with invalidation, our performance gain decreases as the write ratio grows. FarReach [36] is a recent work that enables write-back caching. Although it still cannot serve many workloads due to the size limitation, it provides high performance regardless of the write ratio. The difference between write-through caching and write-back caching is whether a write request for cached items updates the storage server or not. Therefore, StarCache can use write-back caching as well by letting the switch return write replies upon receiving write requests after updating the cache only. However, the switch with Intel Tofino 1 does not have sufficient stages to implement other key modules like snapshot generation alongside the StarCache modules. We expect that we may fully implement write-back caching in StarCache using Tofino 2 [2] whose the number of stages is much higher than Tofino 1 by roughly double, but this is beyond the scope of this work.
@@
Multi-pipeline deployment. The multi-terabits performance of the programmable switch ASIC comes from the multi-pipelined architecture. Each pipeline consists of tens of regular ports and one internal recirculation port. In a switch with 4 pipelines and 64 front ports, each pipeline has 16 ports and one recirculation port statically. The switch resources are partitioned, and metadata and memory data are not shared between the pipelines. Therefore, the pipelines of client-directed ports and server-directed ports should be the same. Otherwise, cache packets in a pipeline may not read requests since they are in another pipeline. This can be achieved by mapping each pipeline of the ToR switch to a client-directed port on the spine switches, which is feasible only when the clients and servers are located in different racks. We expect this to be mitigated with a new programmable switch architecture like MP5 [37] that achieves high performance with a logical single pipeline.
We describe how StarCache can ensure cache coherence between switch pipelines. First, the switch forwards write requests to the recirculation ports of other pipelines and the server. In the other pipelines, the switch drops the request after invalidation as the request has already been delivered to the server in the source pipeline. Meanwhile, write reply packets are cloned in the PRE to fetch the new cache packet for all pipelines. The switch forwards the original reply to the client while recirculating the cloned reply via the recirculation ports of all pipelines. We have implemented this and validated the functionality and performance, but do not include this in the current design. This is because write requests for cached items in a pipeline should be propagated to the other pipelines to invalidate the item value via the recirculation ports, and this may make the port a bottleneck.

### v3

@@ 3.2 Packet Format
Figure 3 depicts the OrbitCache packet format. The message consists of the header and the payload. The switch only parses the header. The payload of OrbitCache message consists of the key and the value. Our header is 22 bytes. Therefore, OrbitCache supports a key-value pair of up to 1438 bytes for a single packet. For example, the switch can cache an item with a 16-byte key and a 1422-byte value. Our header fields are as follows.
@@
Figure 4: Request processing. (a) the switch drops the request after inserting request metadata into the request table; (b) If a circulating cache packet reads request metadata, the switch clones the packet so that the original packet is forwarded to the client and the cloned one is recirculated again for further serving; (c) the switch invalidates the item to avoid inconsistent reads if a write request is for a cached item; (d) upon receiving a write reply for a cached item, the switch validates the item. After that, the switch clones the packet. The cloned packet is processed as a read reply after updating the operation type.
@@
  • FLAG (1 byte): a flag field to distinguish write requests for cached items from those for uncached items.
@@ 3.3 Basic Request and Reply Processing
Request generation at clients. Clients should specify the OP and HKEY fields. The SEQ field is increased by one for every request, which is used to resolve hash collisions. The destination storage server is determined by hashing the key.
@@
Read replies. When receiving a read reply, the switch first checks to see if the ingress port is the recirculation port. If it is, the reply is a cache packet. Otherwise, it is a reply for an uncached item sent by the server. For cache packets, if the cache misses or the state of the value is invalid, the switch drops the packet. A cache miss for a cache packet means that the controller evicted the key from the lookup table due to a change in key popularity or the cache size. The invalid state indicates a write request with a new value is in progress.
The switch goes through the lookup and state tables, and then looks for pending requests in the request table. If a request is found, the switch forwards the cache packet to the client after updating the header with metadata and removing the metadata from the table slot. To make the cache packet serve more requests, the switch clones the packet before forwarding. The switch forwards the original packet to the client and the cloned one to the recirculation port. The switch recirculates the packet if there are no pending requests.
Write requests. The switch checks whether the requested key is in the cache lookup table. If it is, the value for the key is invalidated to prevent reading the outdated value. In addition, the switch sets the FLAG field to 1 to indicate that this request is for a cached item. This makes the storage server append the value in the write reply. Regardless of a cache hit, the switch forwards the request to the storage server to update the value of the key in the server.
Write replies. If the key is cached, the switch validates the value to allow read requests to get the latest value. The switch clones the packet so that the client receives the write reply while the switch has a new cache packet simultaneously. The switch updates the OP field to R-REP of the cloned packet after the first recirculation since cache packets should be read replies. For a cache miss, the switch forwards the packet to the client since the reply is for an uncached item.
Other types of messages. Fetch requests and correction requests are delivered to the server as standard packets. The fetch reply is processed with the same logic as writing replies.
@@ 3.4 Buffering Request Metadata
Circular queue-based request table. We should store multiple request metadata for cached items because the requests should wait until cache packets serve them. The switch should be able to buffer concurrent read requests for the same cached item. If not, many requests for the cached item would be forwarded to the storage server due to the lack of vacancies. Furthermore, a request for a key should be isolated from requests for other keys since the request may undergo excessive queueing delay due to the different requests.
We design a request table based on a circular queue structure. Our request table provides a logical queue for each cached key, and the queue can be accessed in $O(1)$ by efficient indexing. This provides fast and isolated queue access for a key regardless of other keys. The request table consists of 6 register arrays: the first three arrays are for storing request metadata for each key (i.e., client IP address, request sequence number array, and L4 port number). The other three arrays are for managing queue operations. The queue length array maintains the number of stored request metadata for each key. The front pointer array handles dequeue operations by tracking the first request metadata for each key. The rear pointer arrays perform enqueue operations by keeping track of the last request metadata for each key.
The queue management arrays are indexed using a table index ($CacheIdx$) returned by the cache lookup table. The request metadata arrays are accessed via a request index, which is defined as $ReqIdx=CacheIdx\times S+i$ where $S$ is the maximum queue size per key and $i$ is the offset for $i$th slot in the logical queue for a key and is given by the pointer arrays. The switch uses three match-action stages for a request table. Specifically, the switch first checks the queue status (Stage 1), and performs en/dequeue operations if the queue is not full/empty (Stage 2). Next, the switch gets/puts metadata from/into the request table (Stage 3).
@@ 3.5 Cache Serving
Handling multiple requests via packet cloning. In OrbitCache, cache packets pass through the switch data plane repeatably to serve pending requests for cached items. One challenge is serving multiple requests with a cache packet fetched only once because the cache packet is forwarded to the client after updating the header using the retrieved metadata. A strawman is to fetch the cache packet from the server again, but this approach is inefficient as the switch cannot serve pending requests for the key until the fetching is completed. Fetching multiple copies of the cache packet may be a solution, but this incurs excessive queueing delay between cache packets. In addition, it is hard to know how many cache packets should be fetched because we cannot predict the number of requests for each cached item in advance.
@@
To address this, we utilize packet cloning, another built-in feature of the programmable switch ASIC besides packet recirculation. Packet cloning is done by the PRE, a hardware module specialized for packet cloning in the switch ASIC. Cloning has low overhead for the following reasons. First, the PRE is located after the ingress pipeline. This means the switch does not repeat the ingress pipeline processing for the cloned packet, not causing extra ingress processing delay. Second, the switch does not copy the entire packet. It only copies the small descriptor pointing to the memory location of the packet and reuses the packet data.
@@ 3.6 Handling Hash Collisions
Client-side collision resolution. We handle hash collisions at the client level by maintaining a list of the keys for each request that has not yet received a reply. The list is indexed by $pkt.seq$, a request ID. Request packets contain both the original key and the key hash since we should use the original key to get the value in servers. Upon receiving read replies, the client checks whether the requested key in the list and the returned key in the reply header are identical. If different, the client sends a correction request to the storage server. The switch then forwards the request without applying the cache logic so that the client gets the correct item value from the storage server. One issue is that requests for uncached but hash-collided keys always undergo this process. However, this is uncommon since our 128-bit key hash provides a low collision probability. In our experience, hash collisions occasionally occur during cache updates.
An example of resolving hash collisions. We plot an example of handling hash collisions in Figure 7 where the requested key is DDDD but the returned key is AAAA. Upon receiving the reply, the client detects that the retrieved value is a wrong value by referring to the key list using $pkt.seq=1$. The client then sends a correction request to the storage server. The switch bypasses the cache logic, and forwards the packet to the server. The storage server returns the value as a read reply, and the client finally gets the correct value for the key DDDD. The client removes the key from the list. $pkt.seq$ wraps around if it reaches the maximum value.
@@ 3.7 Cache Coherence
To ensure cache coherence between the switch and servers, we design an invalidation-based coherence protocol illustrated in Figure 4 (c) and (d). The switch invalidates the value when handling a write request for a cached item and revalidates the value upon receiving a write reply. Ideally, the outdated cache packets should be replaced with the latest cache packet. However, requests may be returned with the stale value since the outdated packets are still circulating in the switch data plane even if the item value becomes invalid. To address this, the switch drops the cache packet if the item is cached but its value is invalid. Owing to the fact that the cache packet is dropped before accessing the request table, we can prevent read requests from retrieving the stale value until a new cache packet is fetched. The storage server sends a single reply packet, and the switch updates the value and replies to the client simultaneously by cloning the packet.
@@ 3.8 Cache Updates
OrbitCache can handle dynamic workloads by tracking workload changes as follows. To handle key popularity changes over time, we update cache entries based on periodic popularity reports from the switch and storage servers. In particular, we count the popularity of cached keys in the switch data plane using the key popularity counter. The controller keeps track of key popularity by reading the counter periodically. Meanwhile, storage servers periodically report the top-$k$ keys to the controller, which are popular uncached keys. The servers use a count-min sketch with five hash functions to track key popularity in a memory-efficient manner while ensuring accuracy. To reflect the recent status only, we reset all the counters to zero after reporting.
If the cache needs to be updated due to a change in key popularity, the controller deletes the victim key from the cache lookup table and inserts a new popular key instead. The controller then sends a fetch request to the storage server containing the latest value for the key. Upon receiving the request, the server fetches a new cache packet to the switch data plane and replies to the controller. The new popular key inherits the table index ($CacheIdx$) of the evicted key. With this, the pending requests for the evicted key can be handled by the new cache packet and the hash collision resolution mechanism. Note that the hash collision here is due to the same table index, not that the hashes actually collide.
@@ 3.9 Handling Practical Requirements
Failure handling. For packet loss, we can use an application-level mechanism. Our controller uses UDP with a timeout-based mechanism to exchange fetch requests/replies. Storage servers use TCP to report top-$k$ items to the controller. Meanwhile, server failures do not cause a OrbitCache-specific problem. Switch failures result in the loss of cached items, but the cache can be reconstructed quickly by the controller after the switch is recovered because the switch failure is similar to the rapid key popularity changes.
Multi-rack deployment. OrbitCache can be deployed for clusters where multiple racks exist because the ToR switch caches hot items of storage servers belonging to its rack only. For example, assume that we have two racks where ToR switches, $ToR1$ and $ToR2$, are interconnected via a spine switch $SPN$. When $CLI$, a client in rack 1, sends a request to $SRV$, a server in rack 2, the path for an uncached item is $CLI-ToR1-SPN-ToR2-SRV-ToR2-SPN-ToR1-CLI$. If the requested item is cached, $ToR2$ is the only switch that applies the cache logic. There is no issue even if rack 1 is the replicated rack of rack 2 where the two ToR switches have the same cache entries because the request is handled by $ToR1$ directly. If the client is located in another rack and the other two racks are replicated ones, we can balance requests for those racks by implementing load-balancing mechanisms in the ToR switch of the rack where the client exists.
@@ 3.10 Discussion
Dynamic cache sizing. The value size of hot items may change over time, or the value size of a new hot item may differ from that of the evicted item. In this case, the optimal cache size can be changed. To address this, the controller may change the cache size to maximize the performance. The controller calculates the request overflow ratio periodically, which is defined as the number of requests for cached items that were sent to the server divided by the total number of requests for cached items. The controller compares the ratio with a pre-defined threshold (e.g., 1%). If the ratio exceeds the threshold, the controller decreases the cache size. Otherwise, the cache size is increased. We can also set the minimum and maximum cache sizes to prevent overreaction.
Multi-packet items. Many key-value items are less than the MTU size. Therefore, existing works in key-value stores generally consider item values up to 1024 bytes [29, 30, 31, 17, 6, 26]. However, some items may exceed the MTU size, such as articles and photo objects for some workloads (e.g., Wikipedia and Flickr [8]). To cache multi-packet items, we should fetch multiple cache packets with fragmented values for the same key. To do this, we maintain the number of forwarded cache packets for each item by placing another register array alongside the request table, the ACKed packet counter. The initial value of each slot is 1 since most items are single-packet. When fetching item values, the storage server puts the number of packets comprising the item in the FLAG field. When handling cache packets, the switch retrieves metadata as usual, but it does not manipulate the slots of all register arrays in the request table if the slot value in the ACKed packet counter is not equal to the value in FLAG. Instead, the switch increases the slot value by one. When the slot value in the counter equals FLAG, the switch removes request metadata as it is ready to be finished.
However, this may reduce the performance gains since multiple cache packets occupy the cache for one large item. Therefore, the balancing efficiency may decrease if many items consist of multiple packets. One alternative direction is designing a hierarchical cache architecture. We may integrate OrbitCache with Pegasus [27] where the switch replicates hot items across storage servers. Specifically, we can use the switch as the primary cache and the storage servers (or additional cache servers) as secondary caches. For example, we can cache hot single-packet items in the switch while caching hot multi-packet items to servers.
@@
Write-back caching. Although most workloads are read-intensive, some workloads are write-intensive. Since we use write-through caching that updates the switch and the server simultaneously, our performance gain decreases as the write ratio grows. FarReach [34] is a recent work that enables write-back caching. Although it still cannot serve many workloads due to the size limitation, it provides high performance regardless of the write ratio. The difference between write-through caching and write-back caching is whether a write request for cached items updates the storage server or not. Therefore, OrbitCache can use write-back caching as well by letting the switch return write replies upon receiving write requests after updating the cache only, though we need extra modules like snapshot generation.
@@
Multi-pipeline deployment. The multi-terabits performance of the programmable switch ASIC comes from the multi-pipelined architecture. Each pipeline consists of tens of regular ports and one internal recirculation port. The switch resources are partitioned, and metadata and memory data are not shared between the pipelines. Therefore, the pipelines of client-directed ports and server-directed ports should be the same. Otherwise, cache packets in a pipeline may not read requests since they are in another pipeline. This can be addressed by mapping each pipeline of the ToR switch to a client-directed port on the spine switches. This is especially feasible when the clients and servers are located in different racks. We expect this to be addressed more easily with a new programmable switch architecture like MP5 [35] that achieves high performance with a logical single pipeline.